# Role of flow topology in wind-driven wildfire propagation

### v1

@@ Appendix A Implicit-Explicit time integration algorithm
$\hskip 4.26773pt\boldsymbol{\widehat{T}}^{(i)}\leftarrow\boldsymbol{T}^{n}+\Delta\tau\sum_{j=1}^{i-1}\tilde{a}_{ij}F_{j}^{T},\quad\overline{\boldsymbol{T}}^{(i)}\leftarrow\boldsymbol{T}^{n}+\Delta\tau\sum_{j=1}^{i-1}a_{ij}F_{j}^{T},\quad\boldsymbol{\beta}^{(i)}\leftarrow\boldsymbol{\beta}^{n}+\Delta\tau\sum_{j=1}^{i-1}a_{ij}F_{j}^{\beta}$
  compute
$\hskip 9.95863pt\qquad F_{i}^{\beta}\leftarrow-\frac{\epsilon}{q}\frac{\boldsymbol{\beta}^{(i)}s(\widehat{\boldsymbol{T}}^{(i)})^{+}\mathrm{e}^{\widehat{\boldsymbol{T}}^{(i)}/(1+\epsilon\widehat{\boldsymbol{T}}^{(i)})}}{[1+\left(\epsilon/q\right)\Delta ta_{ii}s(\widehat{\boldsymbol{T}}^{(i)})^{+}\mathrm{e}^{\widehat{\boldsymbol{T}}^{(i)}/(1+\epsilon\widehat{\boldsymbol{T}}^{(i)})}]}$
  solve for $F_{i}^{T}$ the linear system:
$\hskip 9.95863pt\qquad F_{i}^{T}=\frac{1}{Da}\nabla\cdot[K(\widehat{\boldsymbol{T}}^{(i)})\nabla(\overline{\boldsymbol{T}}^{(i)}+\Delta\tau a_{ii}F_{i}^{T})]+s(\widehat{\boldsymbol{T}}^{(i)})^{+}[\beta^{(i)}+\Delta\tau a_{ii}F_{i}^{\beta}]\mathrm{e}^{\widehat{\boldsymbol{T}}^{(i)}/(1+\epsilon\widehat{\boldsymbol{T}}^{(i)})}$
$\hskip 99.58464pt-\alpha[\overline{\boldsymbol{T}}^{(i)}+\Delta\tau a_{ii}F_{i}^{T}]-\frac{\overrightarrow{\mathbf{w}}}{\Phi}\cdot\nabla\widehat{\boldsymbol{T}}^{(i)}+\Delta^{2}_{T}(\widehat{\boldsymbol{T}}^{(i)})$
end do
$\hskip 28.45274pt\boldsymbol{T}^{n+1}\leftarrow\boldsymbol{T}^{n}+\Delta\tau\sum_{j=1}^{s}b_{j}F_{j}^{T},\quad\boldsymbol{\beta}^{n+1}\leftarrow\boldsymbol{\beta}^{n}+\Delta\tau\sum_{j=1}^{s}b_{j}F_{j}^{\beta}$
         Marched solution: $\boldsymbol{T}^{n+1}$ and $\boldsymbol{\beta}^{n+1}$ at $\tau^{n+1}=\tau^{n}+\Delta\tau$
Here, $a_{ij}$ and $\tilde{a}_{ij}$ represent the coefficients in a standard Butcher tableau, used to calculate internal temporal stages performing simultaneous explicit and implicit time integration in an IMEX-RK scheme setting.
Appendix B Numerical solver validation
In this section, we validate our solver using a benchmark 2D perfect initial Heaviside firefront over a uniform fuel distribution scenario, as presented in Bürger et al. (2020). Resolving the advecting Heaviside firefront over a long integration time poses a significant numerical challenge, as it excites all possible frequencies, including spurious ones associated with Gibbs’ oscillation in a Fourier spectrum. Despite this challenge, the OUCS2 compact scheme adeptly captures the spatiotemporal evolution of the Heaviside firefront effectively, even without incorporating the LAD model for the considered wildfire scenario (Fig. 9). We maintain consistency by adhering to the parameters in Bürger et al. (2020): a square domain size of $l_{0}=50$, $\epsilon=3\times 10^{-2}$, $q=1$, $\alpha=10^{-3}$, $\overline{T}_{\text{pc}}=3$, initial firefront temperature $\overline{T}=31$, wind speed $|\overrightarrow{\mathbf{w}}|=425$ with a direction of $45^{\circ}$, $C_{\mu}=0$, and $Da=\Phi=1$. Furthermore, with the initial firefront precisely positioned at coordinates $(8.0,8.0)$ and a square side length of 4, and the fuel uniformly initialized at 0.6 across the spatial domain, the simulation is conducted with a time step of $\Delta\tau=10^{-7}$ over a grid size of $256\times 256$. The results depicted in Fig. 9, illustrate the spatio-temporal evolution of the Heaviside firefront at different time instants, demonstrating good agreement with the Bürger et al. (2020) results. Discrepancies may arise from using a more accurate IMEX-RK3 time integration and OUCS2 schemes in our work, contrasting with the employment of IMEX-RK2 scheme and highly dissipative WENO scheme in Bürger et al. (2020).
Appendix C Notations
@@
         $k$          thermal conductivity of fluid
         $C$          specific heat of fluid
@@
         $\overline{R}_{L},\phi_{L}$          Real and phase angle of TF for a left firefront

### v2

@@ Appendix C Methodological Comparison of Wildfire Models
[…]
The results, shown in Fig. 10a, depict the spatio-temporal temperature evolution of the fireline at selected time instants, while Fig. 10b presents the quantitative comparison of the rate of spread ($R_{s}$) across the three wildfire modeling approaches. The CDR model’s predictions are compared against data from Pimont et al. (2012), highlighting both agreements and some discrepancies. The deviations were relatively minor (particularly at lower velocities) and can be attributed to several modeling assumptions, including the use of localized artificial diffusion, spatial parameter averaging, and the omission of complex flame–wind coupling effects. Moreover, such discrepancies are expected due to the inherent limitations and simplifications of the CDR framework. Prieto et al. (2015) presented a two-dimensional simplified CDR wildfire model and demonstrated that the governing spatially averaged physical variables can be optimized through a constrained minimization approach using data from either three-dimensional CFD simulations or experimental measurements, thereby enabling the two-dimensional model to approach more realistic results.
Appendix D Notations
@@
         $k$          Thermal conductivity
         $C$          Specific heat
@@
         $R_{s}$          Rate of Spread
         $R_{0}$          Rate of Spread at zero-slope and no wind condition